OpenStudy (anonymous):

please help, algebra!!! in the u.s., the number Y (in millions) of bicycles sold with wheels sizes of 20in. or greater can be modeled by the equation of Y=0.3X+15 where X is the number of years since 1981. what year were about 22 million bicycles sold? explian

OpenStudy (anonymous):

isn't it hard

OpenStudy (compassionate):

What exactly are you having trouble understanding?

OpenStudy (anonymous):

I don't get how you find the answer\

OpenStudy (compassionate):

Plug in the number 22mil

OpenStudy (anonymous):

where in the Y?

OpenStudy (compassionate):

@Luigi0210 explain this. I'm busy.

OpenStudy (anonymous):

22MIL=0.3X+15like that

OpenStudy (compassionate):

No.

OpenStudy (anonymous):

than where should I put the 22mil?

OpenStudy (compassionate):

You need to do trial and error. We know that it's from 1981 and up. So put various dates, like 2011 or 1998, into the x position.

OpenStudy (anonymous):

so Y=.3 times 2011+ 13

OpenStudy (compassionate):

Mhm

OpenStudy (compassionate):

+ 15

OpenStudy (anonymous):

ya 15

OpenStudy (compassionate):

So you just keep plugging in dates until you reach 22mil. It's trial and error.

OpenStudy (anonymous):

on 2011 I got 618.3

OpenStudy (compassionate):

So plug in 2066 or another number until you get 22mil

OpenStudy (compassionate):

Y=0.3X+15 Imma plug in 5000 y = 0.3(5000) + 15 y = 1515 So plug in something else. y = 0.3(250,000) + 15 y = 75015

OpenStudy (compassionate):

I think this is too complex and you're not getting it. I'll just do it for you.

OpenStudy (compassionate):

22 = 0.3X+15 0.3X = 7 X = 23.3 year = 1981+23 = 2004 Happy?

OpenStudy (compassionate):

Since X is the number of years since 1981 we first disregard the fractional part of X and just add 23 to 1981 because if you are 1/3 through a certain year you are still in that year. Just like if I told you I was 18 even though my birthday was 6 months ago. 1981+23=2004.

OpenStudy (anonymous):

omg thanks so much love realy helped

OpenStudy (anonymous):

@Compassionate